# Learning finitely correlated states: stability of the spectral reconstruction

### v2

@@ Abstract
We show that marginals of blocks of $t$ systems of any finitely correlated translation invariant state on a chain can be learned, in trace distance, with $O(t^{2})$ copies – with an explicit dependence on local dimension, memory dimension and spectral properties of a certain map constructed from the state – and computational complexity polynomial in $t$. The algorithm requires only the estimation of a marginal of a controlled size, in the worst case bounded by the minimum bond dimension, from which it reconstructs a translation invariant matrix product operator. In the analysis, a central role is played by the theory of operator systems. A refined error bound can be proven for $C^{*}$-finitely correlated states, which have an operational interpretation in terms of sequential quantum channels applied to the memory system. We can also obtain an analogous error bound for a class of matrix product density operators reconstructible by local marginals. In this case, a linear number of marginals must be estimated, obtaining a sample complexity of $\tilde{O}(t^{3})$. The learning algorithm also works for states that are only close to a finitely correlated state, with the potential of providing competitive algorithms for other interesting families of states.
@@ 1 Introduction
Obtaining an approximate description of the state of a quantum system consisting of $t$ qudits (also referred to as spins) to decent accuracy without making any assumptions about its structure infamously requires a number of copies of the state that increases exponentially with $t$. This remains true even if these copies are processed by a fully functional quantum computer capable of executing any measurement allowed by quantum mechanics [1, 2]. Consequently, as the number $t$ of spins grows large, the task of learning a description for an arbitrary state of a many-body system, such as a spin chain, becomes exceedingly challenging. A more manageable scenario arises when one is interested in a measurement procedure that ensures approximate learnability under the hypothesis the state belongs to or is close to some specific class of states characterized by a limited number of parameters. In such a case, the number of required copies can be significantly reduced; this is intuitively related to the fact that now a much smaller number of parameters has to be estimated. Tensor network representations provide a means to identify sets of states with controllable expressiveness, quantified by the number of parameters upon which they depend, and remain one of the most widely studied ansatz class. Likewise, ground and thermal states of local Hamiltonians can be fully described by a polynomial number of parameters. In fact, a substantial body of literature has demonstrated the effectiveness of tensor network states for computing or approximating ground and thermal properties of quantum systems. This effectiveness is supported both by theoretical results [3, 4, 5] and in practice [6]. Furthermore, states generated through consecutive applications of a physical map on a hidden memory system also possess the structure of one-dimensional tensor network states. The expressiveness of these states is constrained by the size of the memory system, as illustrated in Figure 1. These are important classes of states on spin chains that can be generated efficiently on typical quantum systems available in laboratories, including quantum processors.
Tensor network states can be considered also on higher-dimensional lattices. In the present work, however, our focus is on the one-dimensional scenario. In this context, tensor network states go by different names, such as matrix product density operators (MPDO), or matrix product states (MPS) in case they are pure. When there are no constraints on positivity, these tensor network representations are referred to as matrix product operators (MPO) or tensor-trains (TT). Specifically, we are interested in certain classes of tensor network states defined on a chain of length $t$, which can be fully reconstructed by knowing the marginals of a fixed size denoted as $t^{*}$. This property is closely related to the concept of injectivity, as discussed in reference [7]. Notably, a special subset of these states corresponds to the marginals of translation invariant states with finite correlations [8], where the knowledge of a marginal of size $t^{*}$ is sufficient for reconstruction.
Our work rests on the crucial fact, elucidated in [8], that finitely correlated states can be seen as a generalization to the quantum setting of stochastic processes admitting finite dimensional linear models, so-called quasi-realizations [9]. In the classical case, a special subclass of the latter go by the name of hidden Markov models [10, 9] (also known as positive realizations). In the quantum case, the natural analogue of hidden Markov models are the aforementioned description of states generated by consecutive applications of a quantum channel to a memory system. We call these models $C^{*}$-realizations. Note however that the class of finite dimensional models is not completely exhausted by finite dimensional $C^{*}$-models, as a recent result by some of the present authors shows [11].
@@
In the classical literature on stochastic processes, learning guarantees for hidden Markov models have been established: a series of works starting with [12] has shown that an estimate of the relevant marginal at precision $\epsilon$ gives rise to a reconstruction of the process which has an error in total variation distance that can be bounded as $O(\operatorname{poly}(t,\epsilon^{-1},m))$ with $t$ being the timespan considered, and $m$ the size of the hidden memory. This in turn implies a rigorous sample complexity bound which scales quadratically in the system size. In the quantum setting, reconstruction algorithms using a similar idea have been proposed, already in the non-translation invariant setting [13, 14] and their practicality and efficiency has been demonstrated with experiments [15]. Moreover, when the reconstruction can be achieved from the knowledge of local marginals of sufficiently small size, local measurements suffice and the quantum complexity of the algorithm is much more tractable than in the general case. However, until the present work, a thorough analysis of the sample complexity of the reconstruction algorithm in the quantum case appeared to be lacking [16]. Here, we precisely propose to fill this gap. Our algorithm and proof strategy build on the classical case, with however important differences due to the presence of entangled observables, and the use of the appropriate operator norms. In fact, our analysis fully exploits the fact that finitely correlated states naturally define an operator system on which one can speak of completely positive maps. This allows us to treat not only the case of states generated by repeated quantum measurements, but also models for which no explanation with finite-dimensional quantum memory exist. Notably, this even provides error bounds for learning classical states that were not previously available. Moreover, the analysis can be generalized to the non-homogeneous case, which is of notable interest since it can address learning of physically motivated states in finite-size systems.
@@ 1.1 Background and related work
As already mentioned, our approach draws inspiration from the literature on hidden Markov models, specifically focusing on what are commonly referred to as “spectral algorithms”. These algorithms are designed to reconstruct a linear model that explains a stochastic process through the estimation of marginals and matrix algebra operations. They aim to overcome the limitations of maximum likelihood approaches, which often lack rigorous convergence guarantees. Spectral algorithms have been applied in more general graphical models as well [17]. The core idea behind these algorithms is to view the probabilities associated with concatenated words as the matrix elements of a mapping between the right segment of the chain and linear functionals on the left segment of the chain. It is assumed that this mapping can be determined by its action on a finite block of the chain. Essentially, these algorithms construct a linear model where the memory system comprises the functionals determined in this manner, and the dynamics are also governed by these functionals. One challenge in ensuring the accuracy of these algorithms lies in their reliance on matrix inversion, which makes them sensitive to small singular values. In a prior study [12], it was demonstrated that the spectral algorithm provides an error bound in terms of total variation distance, assuming that the underlying process closely resembles a hidden Markov model and that certain matrices constructed from hidden Markov model parameters have controlled singular values. This established a learning guarantee akin to PAC (Probably Approximately Correct) learning, which becomes more tractable by narrowing the class of target processes. Subsequent works, among them [18] and [19], relaxed some of the assumptions made in [12], such as the requirement of reconstructing from marginals of size three and invertibility assumptions on the so-called observation matrix (see Section 2.2.1). Nevertheless, a key premise in these works is indeed the existence of an approximating hidden Markov model. To the best of our knowledge, no error bound in total variation distance has been derived for general models (as discussed in [19] and [20]). Notably, there does not even seem to exist an error bound for stochastic processes generated by hidden quantum Markov models, which have been proven to be more expressive than classical ones [21, 22], although they are not as expressive as the most general class of processes one could consider [11]. As a result, the idea of investigating spectral algorithms in the context of hidden quantum Markov models was suggested in [23].
@@
In the quantum information literature, a reconstruction method related to the spectral algorithm we propose is the direct tomography algorithm in [13], which learns a circuit preparing an MPS by knowledge of marginals of sufficiently large size. For general states, the state reconstruction scheme in [14] from the exact knowledge of the marginals, is analogous to the one we present, but is lacking an error bound on the precision of the reconstruction. Moreover, our algorithm uses a slightly different prescription for the reconstruction from empirical data. Note that these algorithms were directly proposed in the non-translation-invariant (non-homogeneous) setting, in contrast to the classical ones. In terms of our error analysis, there is no substantial difference between these two cases except for the fact that, in the non-homogeneous case, one is required to learn several marginal blocks of the state instead of a single arbitrary one. The robustness of the matrix reconstruction in the bipartite case was already discussed in [14] and investigated in more depth in [24], where the authors obtained a bound on the accuracy in terms of the operator norm. However, an error bound in trace distance was not explicitly stated, and it is not clear how to adapt their analysis to get a bound for a chain as opposed to a bipartite state. To perform a similar step in our setting, we instead adapt the analysis carried by [12, 18, 19] to the quantum case. Learning the marginals themselves can be done in several ways, and we take the necessary size of the marginals as a parameter of our class. In the translation invariant case, this size is at most of the order of the minimal dimension of the memory system (bond dimension), while in the non-homogenous case it can be arbitrarily large. That being said, in addition to the use of strictly local measurements to learn the marginals, one can use other methods such as constrained maximum likelihood learning algorithms based on measurement statistics. A very recent result [25] in that direction provides a bound in Hilbert-Schmidt norm for the error in the reconstruction of the state which is polynomial in the size and requires global (in fact Haar random), yet independent, single-copy measurements. It is unclear to us whether this bound directly leads to a good error bound in trace distance. Instead, we use it as a means to learn the marginals with an error in Hilbert-Schmidt distance, which serves our needs adequately as we will show in Section 4.

### v3

@@ Abstract
Matrix product operators allow efficient descriptions (or realizations) of states on a 1D lattice. We consider the task of learning a realization of minimal dimension from copies of an unknown state, such that the resulting operator is close to the density matrix in trace norm. For finitely correlated translation-invariant states on an infinite chain, a realization of minimal dimension can be exactly reconstructed via linear algebra operations from the marginals of a size depending on the representation dimension. We establish a bound on the trace norm error for an algorithm that estimates a candidate realization from estimates of these marginals and outputs a matrix product operator, estimating the state of a chain of arbitrary length $t$. This bound allows us to establish an $O(t^{2})$ upper bound on the sample complexity of the learning task, with an explicit dependence on the site dimension, realization dimension and spectral properties of a certain map constructed from the state. A refined error bound can be proven for $C^{*}$-finitely correlated states, which have an operational interpretation in terms of sequential quantum channels applied to the memory system. We can also obtain an analogous error bound for a class of matrix product density operators on a finite chain reconstructible by local marginals. In this case, a linear number of marginals must be estimated, obtaining a sample complexity of $\tilde{O}(t^{3})$. The learning algorithm also works for states that are sufficiently close to a finitely correlated state, with the potential of providing competitive algorithms for other interesting families of states.
@@ 1 Introduction
Quantum state tomography is a fundamental task that allows the accurate characterization and validation of quantum devices. When provided with a source, such as a quantum circuit or experiment, that generates an unknown quantum system comprising $n$ qudits (or spins), the goal of quantum state tomography is to represent the density matrix of the quantum state through measurements and classical post-processing. It is well established that, in the absence of prior knowledge about the state, attempting to produce an estimate that closely resembles the unknown state requires resources—such as the number of copies, memory or computational time—that increase exponentially with the number of qudits $n$ [1, 2], even when a fully functional quantum computer capable of executing any measurement allowed by quantum mechanics is available. In practice, however, a more manageable situation often arises, as the quantum states of interest typically exhibit a well-defined, physically motivated structure, requiring fewer parameters for full characterization. In these scenarios, one can aim to reconstruct efficient descriptions, with examples of such states including outputs of quantum circuits with bounded gate complexity [3] or thermal states [4]. Alternatively, recent approaches like shadow tomography [5, 6] or classical shadows [7] have emerged, focusing on producing efficient descriptions that yield the same predictions as the density matrix on a subset of observables. In this work, we focus on learning tensor network realizations for states on a one-dimensional lattice, a.k.a Matrix Product Density Operators (MPDOs), with the promise that an efficient realization exists for the otherwise unknown state. When there are no constraints on positivity, these tensor network representations are called simply matrix product operators (MPOs) or tensor-trains (TTs). Translation-invariant states on an infinite chain whose marginals are MPDOs of bounded number of parameters are known as finitely correlated states (FCSs). Tensor network realizations are one of the most widely studied ansatz classes for many-body systems, even beyond the one-dimensional setting we consider. A substantial body of literature has demonstrated the effectiveness of tensor network states for computing or approximating ground and thermal properties of quantum systems. This effectiveness is supported both by theoretical results [8, 9, 10] and in practice [11].
Our work includes the following:
  • A quantum state tomography algorithm for finitely correlated states that can utilize both local and entangled measurements to efficiently reconstruct marginals of the original state using a Matrix Product Density Operator (MPDO) realization (see Fig. 2). Very similar algorithms appeared previously, e.g. in [12].
  • A rigorous sample complexity bound on the trace norm error for these MPDO representations, which scales polynomially with the system size. Our results include both pure and mixed states, providing the first sample complexity guarantee for the tomography of Matrix Product Density Operators from local measurements [13].
  • An extension of our approach to handle the non-translation invariant case, which is particularly relevant for learning physically motivated states in finite-size systems. The learning algorithm is also robust, meaning that it can also be applied to states that are sufficiently close to having an efficient MPDO realization.
Figure 1: Sketch for learning a matrix product operator representation of a finitely correlated state. In the measurement phase we aquire data from $n$ copies of a finite size marginal of the state through local or global measurements. Then the spectral reconstruction phase is a classical post-processing technique that allows us to obtain approximations of the parameters of a linear model that generates the state. Using this approximations we then can generate an accurate estimation of an MPO representation. This allows us to obtain estimations of expected values of local observables or correlations functions on larger marginals of the state.
FCSs were originally introduced to characterize translation-invariant states on quantum spin chains with finite correlations [14]. The initial motivation for these studies was to provide efficient representations of ground states for finite-range interactions, such as the AKLT Hamiltonian [14]. In fact, our approach rests on the crucial observation, elucidated in [14], that finitely correlated states (FCSs) can be viewed as a quantum generalization of stochastic processes which admit finite-dimensional linear models, known as quasi-realizations [15]. At a high level, a quasi-realization of a stochastic process is a collection of matrices and vectors that can reproduce the probabilities of outcomes through simple matrix multiplications and traces. In the classical case, a special subclass of these models is known as hidden Markov models [16, 15] (also referred to as positive realizations). A hidden Markov model is a statistical model where the system being modeled follows a Markov process with unobservable (hidden) states that evolve stochastically, and observable outputs are generated based on measurements of these hidden states. A natural quantum generalization (quantum hidden Markov models) are states generated through consecutive applications of a physical map, which is a quantum channel, on a hidden quantum memory system. We call these models $C^{*}$-realizations, following the nomenclature of [14], where $C^{*}$-finitely correlated states are states admitting such descriptions. This interpretation also motivates the physical interest in learning MPDOs as they include entangled multipartite states that can be generated with a number of local physical operations that scales linearly with the size of the system. Note, however, that the class of finitely correlated states is not completely exhausted by states with finite-dimensional $C^{*}$-models, as a recent result by some of the present authors shows [17] in the case of the infinite chain, while similar statements had been obtained for non-translation invariant states on finite chains [18] and for translation-invariant states on periodic chains [19]. In fact, the most general class of FCSs can be interpreted as being generated via a memory system described by a general probabilistic theory (GPT), which is a mathematical characterization of plausible physical theories that admit state preparation and measurements. The physical interest of states that do not admit finite-dimensional quantum models is an important and unexplored question. In this paper, dealing also with this more general class of states is a natural choice, given the mathematical tools we use.
@@
In the classical literature on stochastic processes, learning guarantees for hidden Markov models have been established for the so-called spectral algorithms: a series of works starting with [20] has shown that an estimate of the relevant marginal at precision $\epsilon^{-1}$ gives rise to a reconstruction of the process with an error in total variation distance that can be bounded as $O(\operatorname{poly}(t,\epsilon^{-1},m))$, where $t$ is the timespan considered, and $m$ is the size of the hidden memory. This, in turn, implies a rigorous sample complexity bound that scales quadratically with the system memory size. In the quantum setting, reconstruction algorithms using a similar idea have been proposed, notably in the non-translation invariant setting [21], and their scalability have been demonstrated with simulations and experiments. In general these approaches involve learning a small marginal of the state using quantum state tomography, reconstructing an approximation of the hidden system, and then using that reconstruction to output a larger marginal of the state. Maximum likelihood methods have also been considered [12]. Alternative methods have been devised for Matrix Product States (MPSs), for which certifiable tomography protocols exist [22] and have been tested in experiments [23]. Moreover, when the reconstruction can be achieved from the knowledge of local marginals of sufficiently small size, local measurements suffice, and the quantum complexity of the algorithm is much more manageable than in the general case.
Our algorithm and proof strategy build on the classical case [20], generalizing the original argument to handle the presence of entangled observables and memory systems beyond classical (and even quantum) systems. At a high level, our algorithm employs what we call the spectral reconstruction technique, in analogy to the classical case. Using this method, given as input the density matrix of marginals of large enough size (say $t^{*}$), using linear algebra (matrix inversion and multiplications), one can reconstruct a valid MPDO representation. In our learning algorithm, we assume that an estimate of the density matrix of the marginal of size $t^{*}$ is provided; errors will naturally occur due to the finite statistics of the experiments or any kind of physical noise; using a modification of the spectral reconstruction, that can better tolerate errors in the estimate, we can then output estimates of model parameters. These can be, in turn, used to construct estimates of marginals at any size $t$, possibly larger than $t^{*}$. While classically this technique was formulated to handle translation-invariant stochastic processes, we also generalize it to address non-translation invariant case. At the technical level, our analysis fully exploits the fact that an FCS naturally defines an operator system [24] on which one can speak of completely positive maps and completely bounded norms, and that the map generating the FCS is contractive. This allows us to address not only states generated by repeated quantum measurements but also models for which no finite-dimensional quantum memory explanation exists. Notably, this even provides error bounds for learning classical states that were not previously available.
@@ 1.1 Background and related work
As already mentioned, our approach draws inspiration from the literature on hidden Markov models, specifically focusing on what are commonly referred to as “spectral algorithms”. These algorithms are designed to reconstruct a linear model that explains a stochastic process through the estimation of marginals and matrix algebra operations. They aim to overcome the limitations of maximum likelihood approaches, which often lack rigorous convergence guarantees. Spectral algorithms have been applied in more general graphical models as well [25]. One challenge in ensuring the accuracy of these algorithms lies in their reliance on matrix inversion, which makes them sensitive to small singular values. In a prior study [20] (see also [26] for a related approach), it was shown that the spectral algorithm provides an error bound in terms of total variation distance, assuming that the process is described by a hidden Markov model satisfying certain assumptions, including that certain matrices constructed from the hidden Markov model parameters have singular values lower bounded by a known constant. This established a learning guarantee akin to PAC (Probably Approximately Correct) learning, which becomes more tractable by narrowing the class of target processes. Subsequent works, among them [27] and [28], relaxed some of the assumptions made in [20], such as the requirement of reconstructing from marginals of size three and invertibility assumptions on the so-called observation matrix. Nevertheless, a key premise in these works is indeed the existence of an approximating hidden Markov model. To the best of our knowledge, no error bound in total variation distance has been derived for general models (as discussed in [28] and [29]). Notably, an error bound for learning stochastic processes generated by hidden quantum Markov models with the spectral algorithm is also missing, which have been proven to be more expressive than classical ones [30, 31], although they are not as expressive as the most general class of processes one could consider [17]. As a result, the idea of investigating spectral algorithms in the context of hidden quantum Markov models was suggested in [32]. We note that for a specific type of quantum hidden Markov models it was shown that it is possible to learn efficiently [33] minimizing the relative entropy between the empirical distribution and distributions in an $\epsilon$-net of the quantum hidden Markov models of fixed memory system. While it is possible that a similar argument could work for a more general class or even for genuinely quantum states, the relative entropy (or any other distinguishability measure) minimization is not computationally efficient, while the spectral algorithms are.
@@
In the quantum information literature, a reconstruction method related to the spectral algorithm we propose is the direct tomography algorithm in [22], which learns a circuit preparing an MPS by knowledge of marginals of sufficiently large size and uses collective operations on subchains of length scaling with the logarithm of the bond dimension. An error guarantee was provided for this algorithm, but it is not clear how to generalize this algorithm for the infinite chain or mixed states. For MPDOs, the state reconstruction scheme in [21] from the exact knowledge of the marginals, is analogous to the one we present, but is lacking an error bound on the precision of the reconstruction. Moreover, our algorithm uses a slightly different prescription for the reconstruction from empirical data. Note that these algorithms were directly proposed in the non-translation-invariant (non-homogeneous) setting, in contrast to the classical ones. In terms of our error analysis, there is no substantial difference between these two cases except for the fact that, in the non-homogeneous case, one is required to learn several marginals of the state instead of a single one. The robustness of the matrix reconstruction in the bipartite case was already discussed in [21] and investigated in more depth in [34], where the authors obtained a bound on the accuracy in terms of the operator norm. However, an error bound in trace distance was not explicitly stated, and it is not clear how to adapt their analysis to get a bound for a chain as opposed to a bipartite state. To perform a similar step in our setting, we instead adapt the analysis carried by [20, 27, 28] to the quantum case. Learning the marginals themselves can be done in several ways, and we take the necessary size of the marginals as a parameter of our class. In the translation invariant case, this size is at most of the order of the minimal dimension of the memory system (bond dimension), while in the non-homogenous case it can be arbitrarily large. That being said, in addition to the use of strictly local measurements to learn the marginals, one can use other methods such as constrained maximum likelihood learning algorithms based on measurement statistics. A very recent result [35] in that direction provides a bound in Hilbert-Schmidt norm for the error in the reconstruction of the state which is polynomial in the size and requires global (in fact Haar random), yet independent, single-copy measurements. It is unclear to us whether this bound directly leads to a good error bound in trace distance. Instead, we use it as a means to learn the marginals with an error in Hilbert-Schmidt distance, which serves our needs adequately as we will show in Section C.